Calculate the speed of sound in air at S.T.P density of air at S.T.P =1.293 kg/m³, y=1.40​

1
Expert's answer
2021-03-21T11:20:42-0400

"v=\sqrt{\frac{\gamma p}{\rho}}=\sqrt{\frac{1.40\cdot 101325}{1.293}}=331.2\ (m/s)" . Answer